# Distributed Quantum Approximate Optimization Algorithm on Integrated High-Performance Computing and Quantum Computing Systems for Large-Scale Optimization

### v1

@@ Abstract
Quantum approximated optimization algorithm (QAOA) has shown promise for solving combinatorial optimization problems by providing quantum speedup on near-term gate-based quantum computing systems. However, QAOA faces challenges in optimizing variational parameters for high-dimensional problems due to the large number of qubits required and the complexity of deep circuits, which limit its scalability for real-world applications. In this study, we propose a distributed QAOA (DQAOA), which leverages a high-performance computing-quantum computing (HPC-QC) integrated system. DQAOA leverages distributed computing strategies to decompose a large job into smaller tasks, which are then processed on the HPC-QC system. The global solution is iteratively updated by aggregating sub-solutions obtained from DQAOA, allowing convergence toward the optimal solution. We demonstrate that DQAOA can handle considerably large-scale optimization problems (e.g., 1,000-bit problem) achieving high accuracy ($\sim$99%) and short time-to-solution ($\sim$276 s). To apply this algorithm to material science, we further develop an active learning algorithm integrated with our DQAOA (AL-DQAOA), which involves machine learning, DQAOA, and active data production in an iterative loop. We successfully optimize photonic structures using AL-DQAOA, indicating that solving real-world optimization problems using gate-based quantum computing is feasible with our strategies. We expect the proposed DQAOA to be applicable to a wide range of optimization problems and AL-DQAOA to find broader applications in material design.
Keywords: HPC-QC integration, large-scale optimization, quantum approximate optimization algorithm, gate-based quantum computing, distributed computing, active learning, material design, high-performance computing
@@ 1.   Introduction
Quantum computing, leveraging quantum phenomena such as superposition and entanglement, significantly advances computational capabilities, promising to solve complex problems that classical computing finds intractable [1, 2, 3, 4, 5, 6]. This transformative technology harnesses the principles of quantum mechanics to process information in fundamentally new ways, revolutionizing various fields including cryptography, materials science, and artificial intelligence [7, 8, 9, 10, 11]. Researchers are increasingly utilizing quantum computing to accelerate their work across various domains [12, 13]. In particular, the quantum approximate optimization algorithm (QAOA) [14], a variational quantum algorithm employing both classical and quantum computing to optimize a cost function by iteratively adjusting parameters in both classical and quantum components (Fig. 1a), has demonstrated significant potential in efficiently tackling complex optimization problems, potentially providing quantum advantage [15, 16, 16, 17]. For example, a recent study by Shaydulin et al. has verified that QAOA can outperform the best-known classical solver for specific optimization problems [18]. Furthermore, other researchers have also applied this quantum algorithm to various combinatorial optimization problems such as the traveling salesman problem, the max-cut problem, and the knapsack problem, demonstrating the potential advantages of leveraging quantum computing [19, 20, 21].
@@
The performance of QAOA has been improved through various strategies, such as advanced ansatz preparation, optimized initialization parameters, and optimized number of layers [22, 23]. However, the current era of quantum computing, known as a noisy intermediate-scale quantum (NISQ) era, presents several challenges. The biggest challenge is that current quantum computing systems, including quantum devices and quantum simulators, are characterized by a limited number of qubits and can only solve problems representing shallow quantum circuits [24, 25, 26]. These limitations make it challenging to scale up to real-world problems (e.g., optimization of photonic structures), which generally require a large number of qubits to represent high-dimensional binary spaces with deep circuits. Consequently, most QAOA studies involve small problems requiring only a few to tens of qubits [27, 19, 28]. For example, Kim et al. demonstrated that QAOA running with Qiskit Aer simulators exhibits low accuracy and long computation times for problems requiring more than 28 qubits, highlighting its difficulty in solving large-scale optimization problems [29].
@@
To address the noise and error challenges in the NISQ era, the research community is exploring hybrid architectures that integrate high-performance computing (HPC) with quantum computing (QC). Here, HPC systems distribute a large task into numerous sub-tasks through distributed computing across multi-cores/nodes, and distributed tasks can be efficiently handled by QC with appropriate sizes. This HPC-QC integrated system can potentially overcome limitations from current quantum systems by distributing heavy workloads and optimizing the use of quantum resources. Consequently, the HPC-QC integrated system can enhance the overall computational capability by leveraging the strengths of both technologies [30, 31, 27]. However, there has been a lack of research on integrating HPC and QC to solve large-scale optimization problems that require hundreds of qubits.
In this work, we develop a distributed QAOA (DQAOA) leveraging an HPC-QC integrated system to efficiently solve real-world optimization problems featuring large-scale models. Our algorithm achieves apparently higher accuracy ($\sim$99%) and is significantly faster (over 160 times) at solving optimization problems (e.g., optimizing metamaterial structures). Furthermore, our approach enables us to efficiently solve large-scale problems (e.g., 1,000-bit problem) while conventional quantum computing approaches find difficulties in solving even much smaller problems (e.g., 30-bit problem). We further propose an active learning algorithm working with the proposed DQAOA for material design. We successfully apply this algorithm to optimize high-performance photonic structures, demonstrating the capability to solve real-world large-scale problems with gate-based quantum computing. By combining the distributed computing capability of HPC with the advantage of QC, our approach aims to tackle the challenges in the NISQ era, paving the way for practical and effective applications of QC in solving complex and large-scale problems. It is important to note that we utilize conventional quantum algorithms and systems, highlighting that the performance of our algorithm can be further improved by employing advanced algorithms and systems. Therefore, our algorithm not only addresses the limitations of current quantum techniques but also sets a foundation for future advancements as quantum technology continues to evolve.
@@ 2.1 Distributed QAOA
Combinatorial optimization problems, known as nondeterministic polynomial-time (NP)-hard problems, can be represented as quadratic unconstrained binary optimization (QUBO) problems, which can serve as Hamiltonians for QAOA [32, 33, 34, 35]. Although QAOA has shown potential for speedup in solving QUBO problems, the limited number of qubits in current quantum systems and the deep circuit depth required for representing large-scale problems (Fig. 1b) present challenges for solving real-world problems in the NISQ era [24, 25, 29]. To address these challenges, we decompose a large QUBO (size of n) into smaller sub-QUBOs (size of k), and solve them through the proposed DQAOA. Here, multi-cores/nodes on HPC systems are employed to solve each sub-QUBO through distributed parallel computing. A global solution $\textit{x}_{g}$ is obtained by aggregating sub-solutions s (from $\textit{s}_{1}$ to $\textit{s}_{b}$, obtained from each sub-task, where b indicates the number of sub-QUBOs). The energy of the large QUBO with a global solution ($\textbf{Q}_{g}$) is calculated, and the ith decision variable of a global solution $\textit{x}_{g,i}$ is updated if $\textbf{Q}_{g}$ decreases by changing the decision variable to that of a sub-solution $\textit{s}_{l,i}$ (i.e., $\textbf{Q}_{g,s_{l,i}}$ < $\textbf{Q}_{g,x_{g,i}}$, Fig. 1c). We first initialize a global solution by solving initial sub-QUBOs that are decomposed with a linear order from the first to the last decision (Fig. S1a). This process ensures that every decision is visited at least once by any sub-QUBO. After that, the original large QUBO is decomposed again into b sub-QUBOs with randomly selected decision variables (according to the sub-QUBO size, k) for DQAOA (Fig. S1b). After solving the sub-QUBOs on different cores and nodes using QAOA, the global solution is updated based on the calculated energy ($\textbf{Q}_{g}$). Additionally, we manually flip the first and last decisions to evaluate the energy of QUBO with the changed binary vector. Iterating these processes improves the quality of the global solution, converging toward a (quasi) optimal solution (Fig. 1c). We note that this algorithm, which integrates classical computing with quantum computing, can work on multi-cores/nodes as well as a single core. QAOA working with a decomposed QUBO on a single core is noted as "dq-QAOA"(decomposed QUBO with QAOA), and it handles a single sub-QUBO without distribution (i.e., the number of sub-QUBO, b in the iteration step, is 1). However, it requires many iterations to achieve a high-quality solution in this case. On the other hand, DQAOA running on multi-cores/nodes HPC systems can solve multiple sub-QUBOs in one iteration through distributed parallel computing (b >> 1), thus requiring significantly fewer iterations to obtain a high-quality solution. For DQAOA, initialization requires at least n-k+1 cores to assign each task (i.e., solving each sub-QUBO) to each core while n is usually much larger than k. Hence, we use more than n cores to solve a problem size of n (i.e., n-bit problem) for DQAOA (see the Method section for more details).
@@
Figure 1: Workflow of DQAOA leveraging an HPC-QC integrated system. a. The schematic of a QAOA circuit, illustrating the iterative process to update variational parameters in both classical and quantum components. b. Circuit depth as a function of a QUBO problem size, showing that the circuit depth increases for larger problems, which indicates the difficulty in solving large-scale problems. c. The schematic of DQAOA to solve large-scale optimization problems through distributed computing. A large QUBO is decomposed into b sub-QUBOs, which are solved by the HPC-QC integrated system.
@@ 2.2 Performance Analysis of DQAOA
We use the Qiskit Aer simulator with the ’automatic’ method for quantum circuits due to its highest accuracy and fastest time-to-solution for QAOA [29], as seen in Fig. S2 (information about evaluating accuracy and time-to-solution is in Supplementary Information). Despite its high performance, the accuracy decreases, and time-to-solution exponentially increases as a problem size (n) increases. Hence, it is hard to solve even 30-bit problems (n=30) using QAOA with this quantum system. It should be noted that we do not propose a new QAOA algorithm or a new quantum simulator to improve their performances. Instead, we propose a strategy that employs the HPC-QC integrated system based on the conventional algorithm and system to scale the problem sizes while achieving high accuracy and reduced time-to-solution. First, we analyze the performance of dq-QAOA for a 30-bit problem, which requires many iterations as it does not leverage distributed computing. Fig. 2a shows that increasing the number of iterations allows for achieving higher accuracy, but the accuracy almost converges after 300 iterations. Notably, more iterations continue to increase time-to-solution, thus the optimal number of iterations is set to 300 for the rest of the studies unless otherwise noted. We further investigate the performance of dq-QAOA with larger sub-QUBO sizes (from 4 to 10). The accuracy does not significantly improve with the larger sizes while time-to-solution exponentially increases due to the increasing number of qubits required and deeper circuits for solving larger sub-QUBOs (Figs. 2b and S2b). Therefore, we use these hyperparameters (the number of iterations of 300 and the sub-QUBO size of 4) for dq-QAOA. As seen in Fig. 2c, QAOA already exhibits a low accuracy for n=30 ($\sim$84%) while dq-QAOA has a much higher accuracy ($\sim$95%). Furthermore, time-to-solution exponentially scales with QAOA while it linearly increases with dq-QAOA, making it several orders of magnitude faster than QAOA (Fig. 2d). Nevertheless, dq-QAOA yields low accuracies for even larger problems ($\geq$50), thus we tune the hyperparameters to verify if higher accuracies can be achieved. Figs. 2e and 2f confirm that increasing the number of iterations and increasing the sub-QUBO size help improve the accuracies for such large problems ($\sim$91$\rightarrow$ $\sim$97for n=50, and $\sim$69$\rightarrow$ $\sim$83for n=150). However, the time-to-solution greatly increases ($\sim$69 s $\rightarrow$ $\sim$2,475 s for n=50, and $\sim$97 s $\rightarrow$ $\sim$2,417 s for n=150), which demands further improvement for the practical application of this algorithm to solve real-world problems.
@@
Figure 2: Performance analysis of dq-QAOA. Accuracy and time-to-solution for a 30-bit problem as a function of (a) the number of iterations, and (b) the sub-QUBO size. The number of iterations is set to 300 and sub-QUBO size to 4 for further studies, as accuracy nearly converges but time-to-solution continues to increase. (c) Accuracy and (d) time-to-solution of QAOA and dq-QAOA as a function of the problem size. Accuracy and time-to-solution of dq-QAOA for a problem size of (e) 50 and (f) 150 with different number of iterations and sub-QUBO sizes. A: 300 iterations with the sub-QUBO size of 4. B: 3,000 iterations with the sub-QUBO size of 4. C: 300 iterations with the sub-QUBO size of 8. D: 1,000 iterations with the sub-QUBO size of 8.
@@
DQAOA performs remarkably better than dq-QAOA and QAOA with significantly higher accuracy and reduced time-to-solution. Fig. 3a indicates that DQAOA achieves a high accuracy ($\sim$95%) and short time-to-solution ($\sim$13 s), needing only 30 iterations, for a large problem (n=150). On the other hand, dq-QAOA shows a lower accuracy ($\sim$83%) with much longer time-to-solution ($\sim$2,417 s) even after the hyperparameter tuning. This notable acceleration in solving the large problem is attributed to distributed parallel computing, which allows the HPC-QC integrated system to handle many sub-QUBOs across multi-cores/nodes simultaneously. Furthermore, this approach enables updating the global solution more times with a larger number of sub-QUBOs, resulting in higher accuracy. Employing even more cores and nodes allows more number of sub-QUBOs (b) for each iteration, possibly leading to obtaining a higher-quality global solution. However, Fig. 3b presents that using cores more than the problem size does not greatly improve the accuracy, but increases time-to-solution due to the communication overheads between cores and nodes. In addition, Fig. S3 reveals that increasing a sub-QUBO size (4 to 8) leads to a slightly higher accuracy but increases time-to-solution quite significantly. Hence, we set the number of iterations to 30, the sub-QUBO size to 4, and the number of cores used equal to the problem size (n). Given these hyperparameters, DQAOA demonstrates a strong capability to scale problem sizes (n=6 to 1,000). DQAOA exhibits a high accuracy ($\sim$99%) and short time-to-solution ($\sim$276 s) even for a very large problem (n=1,000), while QAOA and dq-QAOA cannot handle it (Figs. 3c and 3d). Fig. S4 presents the comparison of accuracy and time-to-solution between QAOA, dq-QAOA, and DQAOA for various problem sizes. The results clearly show that DQAOA exhibits the best accuracy compared to that of QAOA and dq-QAOA (Fig. S4a). Furthermore, DQAOA is >160 times and >7 times faster than QAOA and dq-QAOA, respectively (Fig. S4b). More importantly, DQAOA can further scale problem sizes (over 150), which QAOA and dq-QAOA cannot handle. Note that QUBOs used for this study representing real-world problems are large and dense (Fig. S5), and only DQAOA can solve such large and dense QUBO problems highly efficiently. These performance analysis results demonstrate the superior performance of DQAOA, leveraging the HPC-QC integrated system for distributed computing, in solving large-scale optimization problems, which are intractable with the conventional QAOA. In addition, we expect that the performance of DQAOA can be further improved by employing advanced QAOA algorithms and quantum systems, which is beyond the scope of this study.

### v2

# Distributed Quantum Approximate Optimization Algorithm on a Quantum-Centric Supercomputing Architecture

@@ Abstract
Quantum approximate optimization algorithm (QAOA) has shown promise in solving combinatorial optimization problems by providing quantum speedup on near-term gate-based quantum computing systems. However, QAOA faces challenges for high-dimensional problems due to the large number of qubits required and the complexity of deep circuits, limiting its scalability for real-world applications. In this study, we present a distributed QAOA (DQAOA), which leverages distributed computing strategies to decompose a large computational workload into smaller tasks that require fewer qubits and shallower circuits than necessitated to solve the original problem. These sub-problems are processed using a combination of high-performance and quantum computing resources. The global solution is iteratively updated by aggregating sub-solutions, allowing convergence toward the optimal solution. We demonstrate that DQAOA can handle considerably large-scale optimization problems (e.g., 1,000-bit problem) achieving a high approximation ratio ($\sim$99%) and short time-to-solution ($\sim$276 s), outperforming existing strategies. Furthermore, we realize DQAOA on a quantum-centric supercomputing architecture, paving the way for practical applications of gate-based quantum computers in real-world optimization tasks. To extend DQAOA’s applicability to materials science, we further develop an active learning algorithm integrated with our DQAOA (AL-DQAOA), which involves machine learning, DQAOA, and active data production in an iterative loop. We successfully optimize photonic structures using AL-DQAOA, indicating that solving real-world optimization problems using gate-based quantum computing is feasible. We expect the proposed DQAOA to be applicable to a wide range of optimization problems and AL-DQAOA to find broader applications in material design.
Keywords: Quantum-centric supercomputing, large-scale optimization, quantum approximate optimization algorithm, gate-based quantum computing, distributed computing, active learning, material design
@@ 1. Introduction
Quantum computing, which leverages quantum phenomena such as superposition and entanglement, promises to solve complex problems that appear classically intractable [1, 2, 3, 4]. This transformative technology harnesses the principles of quantum mechanics to process information in fundamentally new ways, revolutionizing various fields including cryptography, materials science, and artificial intelligence [5, 6, 7, 8]. Researchers are increasingly utilizing quantum computing to accelerate their work across various domains [9, 10]. In particular, the quantum approximate optimization algorithm (QAOA) [11], a variational quantum algorithm that leverages both classical and quantum computing to optimize a cost function, has shown particular promise. QAOA iteratively adjusts parameters in both its classical and quantum components (Fig. 1a, see details in the Methods section) and has demonstrated potential for efficiently addressing complex optimization problems [12, 12, 13]. For example, a recent study by Shaydulin et al. has verified that QAOA can outperform the best-known classical solver for specific optimization problems [14]. Furthermore, other researchers have also applied this quantum algorithm to various combinatorial optimization problems such as the traveling salesman problem, Max-Cut problem, and knapsack problem, demonstrating potential advantages of leveraging quantum computing [15, 16, 17].
@@
To improve the performance of QAOA, various strategies are studied, such as advanced ansatz preparation, optimized initial parameters, and optimized number of layers [18, 19]. However, there are several challenges when employing QAOA in practice on today’s quantum devices. The biggest challenge is that current quantum computing systems are characterized by a limited number of qubits and can execute relatively low-depth circuits [20, 21, 22]. These limitations make it challenging to scale up to real-world problems (e.g., optimization of photonic structures), which generally require a large number of qubits to represent high-dimensional binary spaces with deep circuits. Consequently, most QAOA studies involve small problems requiring only a few to tens of qubits [23, 15, 24]. For example, Kim et al. demonstrated that QAOA running with Qiskit Aer simulators exhibits a low approximation ratio and long computation times for problems requiring more than 28 qubits due to an unoptimized number of layers, unoptimized initial parameters, and the barren plateau problem, highlighting its difficulty in solving large-scale optimization problems [25].
@@
The research community has proposed QUBO partition strategies to address such challenges. However, these approaches are primarily focused on sparse problems involving a small number of interactions such as Max-Cut, thus they are less effective when applied to real-world problems characterized by large and dense configurations [26, 27, 15]. Another promising approach is to leverage hybrid architectures that integrate high-performance computing (HPC) with quantum computing (QC). Here, HPC systems distribute a large job into numerous sub-tasks through distributed computing across multi-cores/nodes, leaving the QC to execute circuits amenable to today’s devices. In this way, a quantum-centric supercomputing (QCSC) platform can leverage more effectively current quantum systems by distributing heavy workloads and optimizing the use of limited quantum resources [28]. Consequently, QCSC can enhance the overall computational capability by leveraging the strengths of both technologies [29, 30, 23]. However, there has been a lack of research on integrating HPC and QC to solve large-scale optimization problems, especially those beyond the capabilities of current quantum computers.
In this work, we develop a distributed QAOA (DQAOA) that leverages a QCSC-type architecture to efficiently solve large-scale real-world optimization problems (e.g., optimizing metamaterial structures). Our algorithm achieves a high approximation ratio ($\sim$99%) and exceptional time-to-solution for large-scale problems, outperforming conventional quantum algorithms, including the divide-and-conquer method. Additionally, we execute DQAOA across multiple quantum devices, demonstrating its feasibility for scalable today’s quantum hardware applications. We further propose an active learning algorithm integrated with DQAOA for material design, successfully applying it to optimize high-performance photonic structures. This demonstrates the capability of our approach to address real-world, large-scale problems using current quantum devices. Furthermore, by combining the classical computational power of HPC systems with the advantages of QC, our approach tackles key challenges in materials science, paving the way for practical and effective applications of QC in solving complex optimization tasks.
@@ 2.1 Distributed QAOA
Combinatorial optimization problems, categorized as nondeterministic polynomial-time (NP)-hard problems, can be mapped to quadratic unconstrained binary optimization (QUBO) formulations. These QUBOs can serve as cost Hamiltonians for QAOA [31, 32, 33, 34]. However, translating large and dense QUBOs into QAOA circuits requires a substantial number of two-qubit gates to encode the pairwise interactions between variables, leading to deep circuit depths. Furthermore, the number of qubits needed scales with QUBO sizes, presenting challenges for larger problems. While QAOA has shown promise in solving QUBO problems, its scalability is limited by the number of qubits and the challenges presented by circuit depth, making it difficult to address large-scale, real-world problems on current quantum devices (Fig. 1b) [20, 21, 25]. Note that, as seen in Fig. 1c, a quantum device (IBM-Strasbourg) shows potential acceleration of QAOA jobs compared to a quantum simulator. On the other hand, a quantum emulator (IBM-FakeBrisbane), which simulates quantum processes considering hardware noise via a density matrix, faces significantly higher memory demands, leading to further scaling of time-to-solution compared to a quantum simulator (Fig. S1). Despite the potential advancements of quantum hardware, scalability remains limited due to the excessive number of two-qubit gates. For example, the quantum device faces difficulty in solving QUBO problems larger than 22. A ‘divide-and-conquer’ strategy has been proposed to address scalability challenges, where a large problem is partitioned into several sub-problems that are individually solved using QAOA [26, 27, 15]. While effective for sparse problems, this approach struggles with dense (fully connected) QUBOs, which are common in real-world scenarios [27, 15].
We establish new policies for decomposition and aggregation procedures to efficiently tackle large and dense (fully connected) QUBO problems:
(1) Decomposition policy ensures that correlated interactions between sub-QUBOs in a original dense QUBO are adequately captured. The input QUBO with n variables is decomposed into sub-QUBOs with k variables arranged sequentially from the first to the last decision variable (Fig. S2a). This process guarantees that every decision is visited at least once by any sub-QUBO during the initialization. In the iteration process, k decision variables are randomly selected from the n variables to form sub-QUBOs (Fig. S2b). This process enables the consideration of correlated interactions between different sub-QUBOs. As the number of iterations increases, a broader range of correlated interactions is captured, effectively preserving the critical dependencies in the original QUBO problem.
(2) Aggregation policy ensures that a global solution minimizes a cost function defined by the energy state of the original QUBO. The global solution is initialized during the initialization step by incrementally adding decision variables after solving sub-QUBOs. In the iteration step, each sub-QUBO, generated according to our ‘decomposition policy’, is solved using QAOA, and its sub-solution is used to update the cost. The cost, $\textbf{C}_{g}$, is calculated after every update, which is computationally easy and cheap. The ith decision variable of the global solution $\textit{x}_{g,i}$ is replaced by the corresponding variable $\textit{s}_{l,i}$ from the sub-solution if it results in a reduced cost; i.e., $\textbf{C}_{g,s_{l,i}}$ < $\textbf{C}_{g,x_{g,i}}$, see Fig. 1d. In cases where overlapping variables between sub-QUBOs lead to conflicting decisions, the decision that minimizes the cost is retained. These iterative updates guide the global solution toward the optimal solution by continuously reducing the cost, enabling efficient convergence toward the ground truth of the original QUBO even for dense and large problems.
To accelerate this process, multi-cores/nodes of HPC systems are employed for distributed parallel computing. Each sub-QUBO is assigned to a core or node, and the global solution $\textit{x}_{g}$ is obtained by aggregating sub-solutions s (from $\textit{s}_{1}$ to $\textit{s}_{p}$), where p is the number of sub-QUBOs. After solving all sub-QUBOs, the original large QUBO is decomposed again into p sub-QUBOs with randomly selected decision variables for subsequent iterations (Fig. S2b). Each iteration improves the global solution’s quality, converging toward a (quasi-)optimal solution (Fig. 1d). This distributed quantum-classical algorithm, which we termed Distributed QAOA (“DQAOA"), efficiently integrates quantum computing with HPC resources. It can operate on both multi-core and single-core systems. On a single core, DQAOA simplifies to “dq-QAOA" (decomposed QUBO with QAOA), which handles one sub-QUBO per iteration ($\textit{p}=1$). However, achieving high-quality solutions in this configuration requires many iterations in this case. In contrast, DQAOA leverages distributed parallel computing to solve multiple sub-QUBOs simultaneously ($\textit{p}\gg 1$), significantly reducing the number of iterations needed to reach high-quality solutions. For initialization, DQAOA requires at least $n-k+1$ cores to ensure that each sub-QUBO can be assigned to a separate core. Since n is typically much larger than k, solving an n-bit problem necessitates using more than n cores for DQAOA (see the Method section for more details).
For sparse problems, the sub-QUBOs generated after partitioning a large problem may not have interactions between them (Fig. S3a). In such cases, sub-solutions can be directly aggregated to form a global solution, as overlapping variables between sub-QUBOs do not conflict due to the absence of correlated interactions (Fig. S3b). However, for denser problems, partitioning often results in sub-QUBOs with correlated interactions, complicating the simple aggregation of sub-solutions (Fig. S3c). To address this, our decomposition policy introduces additional sub-QUBOs that capture these correlated interactions (Fig. S3d). The sub-solutions from these additional sub-QUBOs are then used to correct conflicting variables, following our ‘aggregation policy’. Real-world optimization problems, which feature densely (fully) connected interactions, require handling correlations across all sub-QUBOs. To handle these cases effectively, dq-QAOA and DQAOA generate numerous sub-QUBOs, determined by the number of iterations and available cores, ensuring that the global solution progressively improves in quality. This systematic approach effectively captures and preserves the critical dependencies of dense QUBO problems, leading to high-quality global solutions (Fig. S3e).
@@
Figure 1: Workflow of DQAOA leveraging a quantum-centric supercomputing architecture. a, The schematic of a QAOA circuit, illustrating the iterative process to update variational parameters in both classical and quantum components. b, The number of single-qubit and two-qubit gates as a function of QUBO problem size, indicating that circuit depth grows significantly for larger problems, reflecting the increased complexity of solving large-scale optimization tasks. c, Time-to-solution of QAOA for QUBO problems on quantum hardware (IBM-Strasbourg), an emulator (IBM-FakeBrisbane), and a simulator (Qiskit-Aer). The dotted lines indicate the expected time-to-solution for solving large QUBOs (n $>$ 22). d, The schematic of DQAOA to solve large-scale optimization problems through distributed computing. A large QUBO is decomposed into p sub-QUBOs, which are solved by the quantum-centric supercomputing architecture.
@@ 2.2 Performance Analysis of DQAOA
Given the limited access to quantum hardware, we perform the majority of our studies using the Qiskit Aer simulator, which provides near-optimal results for a single QAOA job (Fig. S1) [25]. For performance analysis, we evaluate approximation ratio and time-to-solution metrics (details in Supplementary Information). While QAOA performs well on small problems using the simulator, the approximation ratio decreases and time-to-solution grows exponentially as problem n increases. For instance, solving even a 30-bit problem (n=30) becomes difficult with the simulation. This exponential scaling is primarily due to the quantum simulation being executed on classical computing systems [35]. In quantum simulations, all quantum states must be tracked and stored, leading to an exponential growth in memory requirements as the number of qubits required in quantum circuits increases. This challenge can be effectively mitigated by utilizing quantum hardware, where the quantum state evolution inherently leverages quantum mechanics, avoiding the exponential scaling in the time observed in classical systems (Fig. 1c). It should be noted that this work does not propose a new QAOA or simulator to improve their performances. Instead, we introduce a new strategy based on the conventional algorithm and system to scale the problem sizes while achieving a high approximation ratio and reduced time-to-solution.
First, we analyze the performance of dq-QAOA for a 30-bit problem, which requires many iterations as it does not leverage distributed computing. Fig. 2a shows that increasing the number of iterations allows for achieving a higher approximation ratio, but it approaches convergence after 300 iterations. Notably, more iterations continue to increase time-to-solution, thus the optimal number of iterations is set to 300 for the rest of the studies unless otherwise noted. Next, we investigate the effect of sub-QUBO sizes (ranging from 4 to 10) on dq-QAOA performance. Fig. 2b indicates that while larger sub-QUBO sizes marginally improve the approximation ratio, they exponentially increase time-to-solution due to the characteristics of quantum simulation. For this study, given the use of quantum simulators, we select a sub-QUBO size of 4 for analyzing dq-QAOA.
The advantages of dq-QAOA over standard QAOA are clearly seen in Figs. 2c and 2d. For a 30-bit problem, QAOA achieves a relatively low approximation ratio of $\sim$84%, while dq-QAOA significantly improves it to $\sim$96%. Moreover, the time-to-solution for standard QAOA scales exponentially with the problem size, whereas dq-QAOA demonstrates a linear scaling trend, making it several orders of magnitude faster. Nevertheless, dq-QAOA yields a low approximation ratio for larger problems ($\geq$100). To address this, we tune the hyperparameters by increasing both the number of iterations and the sub-QUBO size. As shown in Figs. 2e and 2f, this hyperparameter tuning improves the approximation ratios for such large problems ($\sim$91$\rightarrow$ $\sim$95for n=100, and $\sim$88$\rightarrow$ $\sim$94for n=150). Since decomposed sub-QUBOs collectively represent the original QUBO, using larger sub-QUBOs and a greater number of sub-QUBOs is more likely to capture essential interactions, leading to improved approximation ratios. However, the time-to-solution significantly increases ($\sim$78 s $\rightarrow$ $\sim$2,134 s for n=100, and $\sim$89 s $\rightarrow$ $\sim$2,405 s for n=150), which highlights the need for further improvement for the practical application to solve real-world problems.
@@
Figure 2: Performance analysis of dq-QAOA. Approximation ratio and time-to-solution for a 30-bit problem as a function of (a) the number of iterations, and (b) the sub-QUBO size. The number of iterations is set to 300 and sub-QUBO size to 4 for further studies, as approximation ratio nearly converges but time-to-solution continues to increase. (c) Approximation ratio and (d) time-to-solution of QAOA and dq-QAOA as a function of the problem size. Approximation ratio and time-to-solution of dq-QAOA for a problem size of (e) 100 and (f) 150 with the different number of iterations and sub-QUBO sizes. A: 300 iterations with the sub-QUBO size of 4. B: 3,000 iterations with the sub-QUBO size of 4. C: 300 iterations with the sub-QUBO size of 8. D: 1,000 iterations with the sub-QUBO size of 8.
@@
DQAOA significantly outperforms dq-QAOA and QAOA with a high approximation ratio ($\sim$95%) and short time-to-solution ($\sim$13 s) for a large problem (n=150) (Fig. 3a). Even after the hyperparameter tuning, for this large-scale problem, dq-QAOA shows a lower approximation ratio ($\sim$94%) and much longer time-to-solution ($\sim$2,405 s). This notable acceleration in solving large problems using DQAOA is attributed to its distributed parallel computing capabilities. By distributing sub-QUBOs across multiple cores/nodes simultaneously, DQAOA updates the global solution more frequently with a larger number of sub-QUBOs, leading to improved approximation ratios. Employing even more cores and nodes allows a greater number of sub-QUBOs (p) for each iteration, potentially leading to a higher-quality global solution. However, Fig. 3b presents that using cores more than the problem size does not greatly improve the approximation ratio, but increases time-to-solution due to the communication overheads between cores and nodes. In addition, Fig. S4 reveals that increasing a sub-QUBO size (4 to 8) leads to a slightly higher approximation ratio but increases time-to-solution quite significantly. Hence, we set the number of iterations to 30, the sub-QUBO size to 4, and the number of cores used equal to the problem size (n). Given these hyperparameters, DQAOA demonstrates a strong capability to scale problem sizes (n=6 to 1,000). Notably, DQAOA achieves a high approximation ratio ($\sim$99%) and short time-to-solution ($\sim$276 s) even for a very large problem (n=1,000), where both QAOA and dq-QAOA cannot handle such scales (Figs. 3c and 3d).